OpenStudy (anonymous):

-5-9 can someone help me

jimthompson5910 (jim_thompson5910):

when you subtract a number from a negative, you can add the two numbers and then make the final answer negative

jimthompson5910 (jim_thompson5910):

so for example -2 - 5 = -7 because 2+5 = 7 then we make the answer negative

jimthompson5910 (jim_thompson5910):

I think of it in terms of debt or IOUs -2-5 means you are 2 dollars in debt, so you have 2 IOUs then you spend 5 more dollars you don't have, so you now have 2+5 = 7 IOUs 7 IOUs = -7 dollars

OpenStudy (anonymous):

thanks for the help

jimthompson5910 (jim_thompson5910):

you're welcome

TheSmartOne (thesmartone):

If you have gotten your answer @bonecheck you can close this question. If you don't know how to do that.. here is a tutrial that will tell you that and more.. http://openstudy.com/study#/updates/543de42fe4b0b3c6e146b5e8

OpenStudy (anonymous):

ok thx

TheSmartOne (thesmartone):

np...

OpenStudy (anonymous):

the smart one do we have to close the problem

TheSmartOne (thesmartone):

Yes if your question has been solved.. I took a screenshot of the part on how to close the question... http://prntscr.com/4xiy3x